Grade 12th passAlgebra

The population of a city decreased from 25,000 to 24,500. Find the percentage decrease.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that
Initial population of the city = 25000
Final population of the city = 24500
Population decrease = Initial population–Final population
= 25000–24500 = 500
Then,
Percentage decrease in population = (population decrease/Initial population) × 100
= (500/25000) × 100
= (50000/25000)
= 50/25
= 2%

​Thanks